OpenStudy (anonymous):

Make the equation have all negative exponents? (x^-4 y^-3) divided by (w^4z) and the entire equation ^5

OpenStudy (anonymous):

To change the sign of an exponent, you change the base to its reciprocal. For example: \[x^{-2}=\frac{1}{x^2}\]\[\left(\frac{x}{y}\right)^{4}=\left(\frac{y}{x}\right)^{-4}\]And so on. Does that help? Or are you still lost?